OpenStudy (anonymous):

Use the function to answer the question. f(x) = -2|x - 5| + 3 What is the vertex of the graph of f(x)?

OpenStudy (anonymous):

vertex of the graph is the point where it shifts from increasing to decreasing or decreasing to increasing ...like two lines meeting at one point...one is -2(x-5)+3 and other line will be -2(-(x-5)) +3 ...equate the two ,yu will get the answer

jigglypuff314 (jigglypuff314):

it is already in vertex form which refers to f(x) = m(x - h) + k m = slope , (h , k) = vertex so for f(x) = -2|x - 5| + 3 ^h ^k

Parth (parthkohli):

The vertex is either the maximum or minimum. The thing here is that if we have a maximum, we won't have a minimum. The graph of these functions with an absolute value is like a "V" extending to infinity to one side. I have more kinds of analyses but trial and test would be OK. -2|x - 5| will be 0 when x = 5 When it's 1, then we have -2|x - 5

Parth (parthkohli):

-2|x+5| would be -2|1+5| = -12

Parth (parthkohli):

Is it -12 for something else too?